The following data relate to marketing expenditure X (in thousand rupees) and the
corresponding sales Y (in ten thousand rupees).
X 10 12 15 20 14
Y 14 15 20 25 10
Find the coefficient of correlation between the marketing expenditure and the corresponding
sales.
1
Expert's answer
2019-04-24T09:57:09-0400

"r = \frac{{}\sum_{i=1}^{n} (x_i - \overline{x})(y_i - \overline{y})}
{\sqrt{\sum_{i=1}^{n} (x_i - \overline{x})^2\sum_{i=1}^{n} (y_i - \overline{y})^2}}=0.7680."